# VELMA: Verbalization Embodiment of LLM Agents for Vision and Language Navigation in Street View

### v1

@@ Abstract
Incremental decision making in real-world environments is one of the most challenging tasks in embodied artificial intelligence. One particularly demanding scenario is Vision and Language Navigation (VLN) which requires visual and natural language understanding as well as spatial and temporal reasoning capabilities. The embodied agent needs to ground its understanding of navigation instructions in observations of a real-world environment like Street View. Despite the impressive results of LLMs in other research areas, it is an ongoing problem of how to best connect them with an interactive visual environment. In this work, we propose VELMA, an embodied LLM agent that uses a verbalization of the trajectory and of visual environment observations as contextual prompt for the next action. Visual information is verbalized by a pipeline that extracts landmarks from the human written navigation instructions and uses CLIP to determine their visibility in the current panorama view. We show that VELMA is able to successfully follow navigation instructions in Street View with only two in-context examples. We further finetune the LLM agent on a few thousand examples and achieve 25%-30% relative improvement in task completion over the previous state-of-the-art for two datasets.
@@ 2 Related Work
Agent models for the outdoor/urban VLN task (Chen et al. 2019) commonly follow a sequence-to-sequence architecture where encoded text and image representations are fused for each decoder step (Xiang, Wang, and Wang 2020; Mehta et al. 2020; Schumann and Riezler 2022). Other proposed agents employ pretrained vision and language transformers that are finetuned on task-specific data (Zhu et al. 2021; Armitage, Impett, and Sennrich 2023). Other work uses CLIP to score the presence of extracted landmarks at each panorama node in a graph and uses this information to plan a route for given navigation instructions (Shah et al. 2022). Their non-urban environment has a graph with 300 nodes, and the navigation path is planned a priori with full access to all panorama images and landmark scores. In contrast, our agent is embodied and has to plan ad-hoc with access to directly observed information only.
@@
Indoor agents (Fried et al. 2018; Wang et al. 2019; Tan, Yu, and Bansal 2019; Fu et al. 2020; Zhu et al. 2020; Qi et al. 2020; Hong et al. 2021; Chen et al. 2021; Li, Tan, and Bansal 2022) are used for navigation datasets like R2R (Anderson et al. 2018) and RxR (Ku et al. 2020). Recently, Zhou, Hong, and Wu (2023) introduced an LLM-based agent for R2R that incorporates image information by transcribing its entire content with an image-to-text model. This is feasible because the navigation trajectories are only six steps on average compared to 40 steps in the urban VLN task considered in our work. Another notable indoor VLN agent uses CLIP to directly predict the next action by scoring the compatibility of the current sub-instruction with available waypoint images (Dorbala et al. 2022).
@@ 3.1 Alignment Inconsistencies in Touchdown
As described in Schumann and Riezler (2022), the automatic rotation mentioned above can lead to generalization problems, e.g., when moving towards the flat side of a T-intersection. For example, if the agent is automatically rotated towards the right facing street and subsequently executes the ${\tt{RIGHT}}$ action, it rotates towards the direction it came from instead of clearing the intersection in the intended direction. The same problem also occurs at intersections with more than three directions. Figure 2 gives an illustrative example that shows the navigation graph at a 4-way intersection. Because the environment is derived from a real-world street layout, the nodes in the graph are not perfectly arranged as in an artificial grid world. In order to make a right turn at the intersection and to follow the route from $v^{1}$ to $v^{5}$, one expects to use the action sequence $[{\tt{FORWARD}},{\tt{FORWARD}},{\tt{RIGHT}},{\tt{FORWARD}},{\tt{FORWARD}}]$. However, when the agent reaches $v^{3}$, it is automatically rotated towards the closest outgoing edge, in this case, $\langle v^{3},v^{4}\rangle$. This is because the rotation $20\degree$$50\degree$ towards $v_{4}$ is shorter than the rotation $20\degree$$345\degree$ towards $v_{7}$. As such, the required sequence of actions to go from $v^{1}$ to $v^{5}$ in Chen et al. (2019)’s environment is $[{\tt{FORWARD}},{\tt{FORWARD}},{\tt{FORWARD}},{\tt{FORWARD}}]$. This is unpredictable and is not correctly aligned with ”turn right at the intersection” instructions. [^1] In Appendix A we show more examples for 3-way, 4-way and 5-way intersections. To alleviate this problem, Schumann and Riezler (2022) explicitly feed the change of heading at each timestep as additional input to their model. This enables the agent to anticipate the unexpected rotation and to adapt to it. Because adding heading delta values to the text-based interface makes it convoluted and unnecessarily difficult for few-shot learning, we propose a more intuitive way to solve this ambiguity at intersections. We modify the state transition function $\phi$ such that the agent is not automatically rotated when moving ${\tt{FORWARD}}$. This means the agent’s heading $\alpha_{t}$ is not automatically aligned with an outgoing edge. Instead, the direction is selected in relation to all outgoing edges. The agent at node $v^{3}$ in Figure 2 has the nodes $v^{6}$, $v^{7}$ and $v^{4}$ in front. The forward direction is selected as the middle one of the three edges, the right direction as the right-most edge, and the left direction as the left-most edge. This means that executing the ${\tt{RIGHT}}$ action at position $v^{3}$ will now rotate the agent towards node $v^{4}$ and allows to use the semantically correct sequence of actions for the depicted route. The proposed modification solves the issue of inconsistent action sequences at intersections and allows to use agents that are not specifically trained in this environment.
@@ 4.2 Datasets
There are two datasets that provide navigation instructions for the environment described in Section 3: Touchdown (Chen et al. 2019) and Map2seq (Schumann and Riezler 2021). Each dataset includes around 10k navigation instances, and we utilize them in the more challenging unseen scenario introduced by Schumann and Riezler (2022). This means that generalization is crucial because the training routes are located in an area that is geographically separated from the area of development and test routes. The main difference between the two datasets is that Touchdown instructions were written by annotators who followed the route in Street View, while Map2seq instructions were written by annotators that saw a map of the route. The Map2seq navigation instructions were subsequently validated to also be correct in Street View. Another difference is that the initial state in Map2seq orientates the agent towards the correct direction which leads to overall better task completion rates than for Touchdown instances.
@@ 5.1 Landmark Extractor
Each navigation instructions text $n$ mentions multiple landmarks as visual guidance. In order to determine if a mentioned landmark is visible in the current panorama view, we first have to extract them from the instructions text. For this, we create a single prompt that includes five in-context examples of navigation instructions paired with a list of landmarks (shown in Appendix B). It is used by the LLM to generate the list of landmarks $(l_{1},l_{2},...,l_{L})$ mentioned in the given navigation instructions. The landmark extractor is depicted in the top middle of Figure 3 and is applied once, before the navigation starts.
@@ 5.2 Landmark Scorer
At each step, the agent observes a panorama view $p^{\alpha}_{v}$, defined by its current position $v$ and heading direction $\alpha$. The view is an 800x460 sized image cut from the panorama with 60° field-of-view. In order to determine if a landmark $l_{i}$ is visible in the view, we employ a CLIP model (Radford et al. 2021) to embed the image and the landmark text: ”picture of [$l_{i}$]”. The similarity score of the two embeddings determines the visibility of the landmark. Because the scores can be biased towards certain types of landmarks, we standardize them using all views $p^{*}_{train}$ of the ~20k panorama images in the training area. Recall that we operate in the unseen scenario where the training area and evaluation area are geographically separated. The standardized score of a landmark is:
@@
If the standardized score is larger than the threshold $\tau$, the landmark is classified as visible in the current view. There are no ground truth labels indicating whether landmarks are visible in certain panoramas, thus the classification is completely unsupervised. The threshold is the only tunable parameter in the landmark scorer. Figure 4 shows the distribution of unstandardized CLIP scores and views at different threshold values for two example landmarks. While the views at $\tau=4.0$ both show the correct landmark, the view at $\tau=3.0$ for ”Bank of America” shows an HSBC branch, and for ”yellow truck” it shows a white truck. This suggests that the optimal threshold lies between the two values. As depicted on the right in Figure 3, the agent also evaluates views to the left and right of the current heading. Each panorama view direction $(p^{\alpha-90\degree}_{v},p^{\alpha-45\degree}_{v},p^{\alpha}_{v},p^{\alpha+45\degree}_{v},p^{\alpha+90\degree}_{v})$ is associated with a string literal $m$ valued left, slightly left, ahead, slightly right or right, respectively. A visible landmark $l_{i}$ and the corresponding direction literal $m_{i}$ are passed to the verbalizer. A full navigation trajectory includes around 200 image views (40 steps and 5 view directions per step) and each landmark is typically visible in only one or two views.
@@ 5.3 Verbalizer
The verbalizer is a template-based component that produces environment observations in text form. There are two types of environment observations. First, there are street intersections that are detected based on the number of outgoing edges $N(v)$ at the current node $v$ in the navigation graph. If there are three or more outgoing edges at step $t$, the verbalizer encodes this information into the observation string $o_{t}^{e}$: ”There is a $[N(v)]$-way intersection”. Extracting this information directly from the navigation graph is akin to the junction type embedding used by the ORAR model (Schumann and Riezler 2022) and motivated by direction arrows displayed in the Street View GUI that human navigators used during data collection. The other type of observations are landmarks visible in the panorama view. The landmark name $l_{i}$ and direction literal $m_{i}$ are used to verbalize the observation $o_{t}^{l}$: ”There is $[l_{i}]$ on your $[m_{i}]$”. The complete observation is $o_{t}=[o_{t}^{e},o_{t}^{l}]$, where the respective string is empty if no intersection or landmark is detected. The observation is appended to the prompt in Equation 1 and used by the agent to decide the next action.
@@ 6 Experiments
We conducted experiments [^2] Project page: https://map2seq.schumann.pub/vln/velma/ and code: https://github.com/raphael-sch/VELMA to evaluate the navigation performance of the proposed LLM agent in finetuning and in-context learning settings. We ran the landmark extractor once for all instances using GPT-3 (Brown et al. 2020) and used the same extracted landmarks in all experiments. Manual inspection shows that GPT-3 reliably extracts all landmarks without obvious errors. We used CLIP-ViT-bigG-14-laion2B-39B-b160k (Schuhmann et al. 2022) as the CLIP model in the landmark scorer. We set the threshold $\tau=3.5$ for all experiments. The threshold was selected by inspecting the distribution of CLIP scores (as in Figure 4) for a handful of landmarks. On purpose, we did not systematically tune it in order to not violate the premise of few-shot learning.
@@ 6.1 Metrics and Baseline
Development Set Test Set
Touchdown Map2seq Touchdown Map2seq
Model SPD↓ KPA↑ TC↑ SPD↓ KPA↑ TC↑ SPD↓ KPA↑ TC↑ SPD↓ KPA↑ TC↑
Seq-to-Seq RNN, full training set
ORAR 20.0 ±0.7 -  15.4 ±2.2 11.9 ±0.4 -  27.6 ±1.8 20.8 ±0.6 -  14.9 ±1.2 13.0 ±0.3 -  30.3 ±1.8
ORAR* 16.5 ±0.1 64.0 ±0.2  22.6 ±0.6 10.3 ±0.4  74.4 ±0.8  29.9 ±1.7 17.4 ±0.2 62.3 ±0.1  19.1 ±1.0 10.9 ±0.1  74.7 ±0.2  32.5 ±1.4
2-Shot In-Context Learning
VELMA-GPT-3 22.2  49.1  6.8 19.1  58.1  9.2 - - - - - -
VELMA-GPT-4 21.8 56.1 10.0 12.8 70.1 23.1 - - - - - -
LLM Finetuning, full training set
VELMA-FT 18.6 ±0.3  61.3 ±0.6  22.6 ±0.4 9.8 ±0.5  76.8 ±0.8  38.2 ±0.5 19.4 ±0.1  60.0 ±0.4  21.0 ±0.2 10.3 ±0.4  76.4 ±0.6  37.8 ±0.9
VELMA-RBL 16.1 ±0.1  63.5 ±0.1 25.7 ±0.4 8.4 ±0.7 78.9 ±0.8 42.8 ±0.6 16.6 ±0.2  61.9 ±0.2 24.1 ±0.7 8.8 ±0.2 77.9 ±0.2 42.1 ±1.1
Table 2: Results for the urban VLN task on Touchdown and Map2seq. ORAR is the previous best model and follows a seq-to-seq architecture that processes text and image features. ORAR* and following models make use of the improved environment introduced in Section 3. The proposed environment and workflow allow to perform the urban VLN task as text-only next word prediction. VELMA-GPT-3 and VELMA-GPT-4 models are prompted with two in-context examples. Due to cost and data leakage concerns, we evaluate the GPT models on the development sets only. VELMA-FT is based on LLaMa-7b that is finetuned on all ~ 6k training text sequences. VELMA-RBL training is described in Section 6.3. All results are for the unseen scenario where evaluation routes are geographically separated from training routes. Experiments are repeated three times with different random seeds (mean/std reported). Bold values are the nominal best results and underlined are best few-shot results.

### v2

@@ Abstract
Incremental decision making in real-world environments is one of the most challenging tasks in embodied artificial intelligence. One particularly demanding scenario is Vision and Language Navigation (VLN) which requires visual and natural language understanding as well as spatial and temporal reasoning capabilities. The embodied agent needs to ground its understanding of navigation instructions in observations of a real-world environment like Street View. Despite the impressive results of LLMs in other research areas, it is an ongoing problem of how to best connect them with an interactive visual environment. In this work, we propose VELMA, an embodied LLM agent that uses a verbalization of the trajectory and of visual environment observations as contextual prompt for the next action. Visual information is verbalized by a pipeline that extracts landmarks from the human written navigation instructions and uses CLIP to determine their visibility in the current panorama view. We show that VELMA is able to successfully follow navigation instructions in Street View with only two in-context examples. We further finetune the LLM agent on a few thousand examples and achieve around 25% relative improvement in task completion over the previous state-of-the-art for two datasets.
@@ 2 Related Work
Agent models for the outdoor/urban VLN task (Chen et al. 2019) commonly follow a sequence-to-sequence architecture where encoded text and image representations are fused for each decoder step (Xiang, Wang, and Wang 2020; Hermann et al. 2020; Mehta et al. 2020; Schumann and Riezler 2022; Sun et al. 2023). Other proposed agents employ pretrained vision and language transformers that are finetuned on task-specific data (Zhu et al. 2021; Armitage, Impett, and Sennrich 2023). Zhong et al. (2021) represent the visual environment by symbols using semantic segmentation and extreme downsampling of panorama images, but their agent does not improve over previous success rates. Other work uses CLIP to score the presence of extracted landmarks at each panorama node in a graph and uses this information to plan a route for given navigation instructions (Shah et al. 2022). Their non-urban environment has a graph with 300 nodes, and the navigation path is planned a priori with full access to all panorama images and landmark scores. In contrast, our agent is embodied and has to plan ad-hoc with access to directly observed information only.
@@
Indoor agents (Fried et al. 2018; Wang et al. 2019; Tan, Yu, and Bansal 2019; Fu et al. 2020; Zhu et al. 2020; Qi et al. 2020; Hong et al. 2021; Chen et al. 2021; Li, Tan, and Bansal 2022) are used for navigation datasets like R2R (Anderson et al. 2018) and RxR (Ku et al. 2020) or ObjectNav (Ramakrishnan et al. 2021; Zhou et al. 2023). Khandelwal et al. (2022) showed that using the CLIP encoder for image features improves performance for a range of vision and language tasks. Recently, Zhou, Hong, and Wu (2023) introduced an LLM-based agent for R2R that incorporates image information by transcribing its entire content with an image-to-text model. This is feasible because the navigation trajectories are only six steps on average compared to 40 steps in the urban VLN task considered in our work. Another notable indoor VLN agent (Dorbala et al. 2022) uses CLIP to directly predict the next action by scoring the compatibility of a sub-instruction with available waypoint images.
@@ 3.1 Alignment Inconsistencies in Touchdown
As described in Schumann and Riezler (2022), the automatic rotation mentioned above can lead to generalization problems, e.g., when moving towards the flat side of a T-intersection. For example, if the agent is automatically rotated towards the right facing street and subsequently executes the ${\tt{RIGHT}}$ action, it rotates towards the direction it came from instead of clearing the intersection in the intended direction. The same problem also occurs at intersections with more than three directions. Figure 2 gives an illustrative example that shows the navigation graph at a 4-way intersection. Because the environment is derived from a real-world street layout, the nodes in the graph are not perfectly arranged as in an artificial grid world. In order to make a right turn at the intersection and to follow the route from $v^{1}$ to $v^{5}$, one expects to use the action sequence $[{\tt{FORWARD}},{\tt{FORWARD}},{\tt{RIGHT}},{\tt{FORWARD}},{\tt{FORWARD}}]$. However, when the agent reaches $v^{3}$, it is automatically rotated towards the closest outgoing edge, in this case, $\langle v^{3},v^{4}\rangle$. This is because the rotation $20\degree$$50\degree$ towards $v_{4}$ is shorter than the rotation $20\degree$$345\degree$ towards $v_{7}$. As such, the required sequence of actions to go from $v^{1}$ to $v^{5}$ in Chen et al. (2019)’s environment is $[{\tt{FORWARD}},{\tt{FORWARD}},{\tt{FORWARD}},{\tt{FORWARD}}]$. This is unpredictable and is not correctly aligned with ”turn right at the intersection” instructions. [^1] In the Appendix we show more examples for 3-way, 4-way and 5-way intersections. To alleviate this problem, Schumann and Riezler (2022) explicitly feed the change of heading at each timestep as additional input to their model. This enables the agent to anticipate the unexpected rotation and to adapt to it. Because adding heading delta values to the text-based interface makes it convoluted and unnecessarily difficult for few-shot learning, we propose a more intuitive way to solve this ambiguity at intersections. We modify the state transition function $\phi$ such that the agent is not automatically rotated when moving ${\tt{FORWARD}}$. This means the agent’s heading $\alpha_{t}$ is not automatically aligned with an outgoing edge. Instead, the direction is selected in relation to all outgoing edges. The agent at node $v^{3}$ in Figure 2 has the nodes $v^{6}$, $v^{7}$ and $v^{4}$ in front. The forward direction is selected as the middle one of the three edges, the right direction as the right-most edge, and the left direction as the left-most edge. This means that executing the ${\tt{RIGHT}}$ action at position $v^{3}$ will now rotate the agent towards node $v^{4}$ and allows to use the semantically correct sequence of actions for the depicted route. The proposed modification solves the issue of inconsistent action sequences at intersections and allows to use agents that are not specifically trained in this environment.
@@ 4.2 Datasets
There are two datasets that provide navigation instructions for the environment described in Section 3: Touchdown (Chen et al. 2019) and Map2seq (Schumann and Riezler 2021). Each dataset includes around 10k navigation instances, and we utilize them in the more challenging unseen scenario introduced by Schumann and Riezler (2022). This means that generalization is crucial because the training routes are located in an area that is geographically separated from the area of development and test routes. The main difference between the two datasets is that Touchdown instructions were written by annotators who followed the route in Street View, while Map2seq instructions were written by annotators that saw a map of the route. The Map2seq navigation instructions were later validated to also be correct in Street View. Another difference is that the initial state in Map2seq orientates the agent towards the correct direction which leads to overall better task completion rates than for Touchdown instances.
@@ 5.1 Landmark Extractor
Each navigation instructions text $n$ mentions multiple landmarks for visual guidance. In order to determine if a mentioned landmark is visible in the current panorama view, we first have to extract them from the instructions text. For this, we create a single prompt that includes five in-context examples of navigation instructions paired with a list of landmarks (shown in the Appendix). It is used by the LLM to automatically generate the list of landmarks $(l_{1},l_{2},...,l_{L})$ mentioned in the given navigation instructions. The landmark extractor is depicted in the top middle of Figure 3 and executed before the navigation starts.
@@ 5.2 Landmark Scorer
At each step, the agent observes a panorama view $p^{\alpha}_{v}$, defined by its current position $v$ and heading direction $\alpha$. The view is an 800x460 sized image cut from the panorama with 60° field-of-view. In order to determine if a landmark $l_{i}$ is visible in the view, we employ a CLIP model (Radford et al. 2021) to embed the image and the caption: ”picture of [$l_{i}$]”. The similarity score of the two embeddings determines the visibility of the landmark. Because the scores can be biased towards certain types of landmarks, we standardize them using all views $p^{*}_{train}$ of the ~20k panorama images in the training area. Recall that we operate in the unseen scenario where the training area and evaluation area are geographically separated. The standardized score of a landmark is:
@@
If the standardized score is larger than the threshold $\tau$, the landmark is classified as visible in the current view. The process does not require annotations and is completely unsupervised, allowing to score novel landmarks. The threshold is the only tunable parameter in the landmark scorer. Figure 4 shows the distribution of unstandardized CLIP scores and views at different threshold values for two example landmarks. While the views at $\tau=4.0$ both show the correct landmark, the view at $\tau=3.0$ for ”Bank of America” shows an HSBC branch, and for ”yellow truck” it shows a white truck. This suggests that the optimal threshold lies between the two values. As depicted on the right in Figure 3, the agent also evaluates views to the left and right of the current heading. Each panorama view direction $(p^{\alpha-90\degree}_{v},p^{\alpha-45\degree}_{v},p^{\alpha}_{v},p^{\alpha+45\degree}_{v},p^{\alpha+90\degree}_{v})$ is associated with a string literal $m$ valued left, slightly left, ahead, slightly right or right, respectively. A visible landmark $l_{i}$ and the corresponding direction literal $m_{i}$ are passed to the verbalizer. A full navigation trajectory includes around 200 image views (40 steps and 5 view directions per step) and each landmark is typically visible in only one or two views.
@@ 5.3 Verbalizer
The verbalizer is a template-based component that produces environment observations in text form. There are two types of environment observations. First, there are street intersections that are detected based on the number of outgoing edges $N(v)$ at the current node $v$ in the navigation graph. If there are three or more outgoing edges at step $t$, the verbalizer encodes this information into the observation string $o_{t}^{e}$: ”There is a $[N(v)]$-way intersection”. Extracting this information directly from the navigation graph is akin to the junction type embedding used by the ORAR model (Schumann and Riezler 2022) and is motivated by direction arrows displayed in the Street View GUI that human navigators used during data collection. The other type of observations are landmarks visible in the panorama view. The landmark name $l_{i}$ and direction literal $m_{i}$ are used to verbalize the observation $o_{t}^{l}$: ”There is $[l_{i}]$ on your $[m_{i}]$”. The complete observation is $o_{t}=[o_{t}^{e},o_{t}^{l}]$, where the respective string is empty if no intersection or landmark is detected. The observation is appended to the prompt in Equation 1 and used by the agent to decide the next action.
@@ 6 Experiments
Development Set Test Set
Touchdown Map2seq Touchdown Map2seq
Model SPD↓ KPA↑ TC↑ SPD↓ KPA↑ TC↑ SPD↓ KPA↑ TC↑ SPD↓ KPA↑ TC↑
ORAR-ResNet 20.0 ±0.7 -  15.4 ±2.2 11.9 ±0.4 -  27.6 ±1.8 20.8 ±0.6 -  14.9 ±1.2 13.0 ±0.3 -  30.3  ±1.8
[…]
2-Shot In-Context Learning
VELMA-Mixtral 28.4  47.2  6.5 21.1  56.8  8.0 - - - - - -
VELMA-GPT-3 22.2  49.1  6.8 19.1  58.1  9.2 - - - - - -
VELMA-GPT-4 21.8 56.1 10.0 12.8 70.1 23.1 - - - - - -
LLM Finetuning, full training set
VELMA-FT 18.3 ±0.4  62.0 ±0.1  23.4 ±0.2 8.7 ±0.0  78.7 ±0.3  41.3 ±0.9 18.2 ±0.3  62.2 ±0.2  23.5 ±0.4 9.7 ±0.3  78.0 ±0.1  40.0 ±1.0
VELMA-RBL 15.5 ±0.3  63.6 ±0.6 26.0 ±0.6 8.3 ±0.1 79.5 ±0.4 45.3 ±0.5 16.0 ±0.7  62.8 ±1.3 26.4 ±1.7 8.3 ±0.2 79.6 ±0.4 47.5 ±0.7
[…]
We conducted experiments [^2] Project page: https://velma.schumann.pub/ and code: https://github.com/raphael-sch/VELMA to evaluate the navigation performance of the proposed LLM agent in finetuning and in-context learning settings. We used CLIP-ViT-bigG-14-laion2B-39B-b160k (Schuhmann et al. 2022) as the CLIP model in the landmark scorer. We set the threshold $\tau=3.5$ for all experiments. The threshold was selected by inspecting the distribution of CLIP scores (as in Figure 4) for a handful of landmarks. On purpose, we did not systematically tune it in order to not violate the premise of few-shot learning.
6.1 Landmark Extraction
We ran the landmark extractor once for all instances using GPT-3 (Brown et al. 2020) and used the same extracted landmarks in all experiments. On average, 2.7 landmarks were extracted from a navigation instructions text. Around 58% of the landmarks in the test sets are novel, i.e., they are not used in the training instances. In order to estimate the quality of the automatically extracted landmarks, we annotated 50 instances of each development set by hand. For Touchdown we calculated an F1-score of 96.3 (precision: 97.2, recall: 95.4) and the F1-score for Map2seq is 99.6 (precision: 100, recall: 99.3). This shows that GPT-3 reliably extracts landmarks from the instructions text and reusing them for all experiments is minimizing the inaccuracies introduced by this workflow step.